OpenStudy (anonymous):

Let f(x)=x^2-16 and g(x)=x+4. Find f/g and its domain.

OpenStudy (anonymous):

@BangkokGarrett help please!!

hero (hero):

\[\dfrac{f}{g}(x) = \dfrac{f(x)}{g(x)} = \dfrac{x^2 - 16}{x + 4}\]

OpenStudy (anonymous):

so it would be x-4..... how can I find a domain for that though?

OpenStudy (anonymous):

Nevermind I got it. Thanks!

hero (hero):

What do you believe is the domain?